OpenStudy (anonymous):

A regular octagon has side length of 10.9 in. The perimeter of the octagon is 87.2 in & the area is 392.4 in sq. A second octagon has side lengths equal to 16.35 in. Find the area of the 2nd octagon. A)882.9 B)717.06 C)642.66 D)461.60

OpenStudy (campbell_st):

the ratio of the side lengths is needed 10.9: 16.35 is the same as 2:3 to compare area square the ratios 2^2:3^2 or 4:9 the area of the larger octagons is 9/4 the smaller area so Larger Octagon Area = 9/4 * 392.4 = 887.4 in^2

OpenStudy (anonymous):

That wasnt one of the choices they gave me.

OpenStudy (campbell_st):

oops my typo... 9/4 * 392.4 = 882.9

OpenStudy (anonymous):

Area of a regular octagon is given by \[ 2(1+\sqrt{2})a^2\] Thus, a regular octagon has side length of 10.9 in will have an area of = \[ 2(1+\sqrt{2})(10.9)^2)= 573.665 \neq 392.4 \]

OpenStudy (campbell_st):

its a question about know the scaling factor for length. and area... not so much about area of a polygon

OpenStudy (anonymous):

The information seems incorrect.

OpenStudy (campbell_st):

yep... lots of those type of things in textbooks.... "not to scale"...lol

OpenStudy (anonymous):

wait so is the answer 882.9